• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 00:18:07 on July27,2006

    Entry number 175483

    keyword=Spectrometers not ready...

    We are not able to contact Tech-on-call or the alternate; quite understandable, given the lateness of today's configuration change (plus the ongoing problems over the past week with spectrometers).

    So, currently the Right Arm Q1 is showing the same fault is displayed earlier in the week when the lead-flow controller went bad ("Earth Leakage Failur" and "Spare Interlock"), so this is potentially the problem.

    The Left Arm Q2 shows a Quench fault, which needs to be reset (cooling flows all seem ok).

    If/when beam comes, we will proceed with non-spectrometer tests: beam-size checks, raster settings, maybe radiator commissioning, BPM calibration, BCM re-calibration. If the beam doesn't come for a couple of hours, this will likely take us until morning when Techs will be in to reset the quads. Should be all ok.